# MHE under parametric uncertainty - Robust state estimation without informative data

### v1

@@ 7 Conclusion
State estimation algorithms require some form of online adaptation to address nonlinear systems with unknown or time-varying parameters. In this paper, we showed that standard moving horizon estimation (MHE) for joint state and parameter estimation can fail in the absence of persistency of excitation (PE), e.g., if a system is well regulated or over-parametrized. We proposed a novel MHE formulation with regularization on a (constant) prior estimate of the unknown system parameter and established practical stability of the resulting state estimate independent of PE. We discussed how many existing results from MHE, adaptive estimation, and functional estimation relate to the proposed MHE approach. Finally, we showed practical applicability using a car example.
While the proposed regularization on a given prior parameter estimate allows us to establish stability irrespective of PE, it may deteriorate the estimation performance compared to a standard MHE approach in case PE is sufficiently large. An interesting question for future research is to extend the proposed MHE approach to achieve an estimation performance comparable to a standard MHE scheme in case of large PE, while the robust stability properties irrespective of PE of the proposed approach are preserved.

### v2

@@ 7 Conclusion
State estimation algorithms require some form of online adaptation to address nonlinear systems with unknown or time-varying parameters. In this paper, we showed that standard moving horizon estimation (MHE) approaches for joint state and parameter estimation can fail in the absence of persistency of excitation (PE), e.g., if a system is well regulated or over-parametrized. To overcome this issue, we proposed the use of a regularization based on a (constant) prior estimate of the unknown parameter within an MHE formulation and established practical stability of the resulting state estimate independent of PE for general nonlinear systems. We discussed how many existing results from MHE and adaptive estimation relate to the proposed MHE approach. Finally, we showed practical applicability using a car example.
While the proposed regularization on a given prior parameter estimate allows us to establish stability irrespective of PE, it may deteriorate the estimation performance compared to a standard MHE approach for joint state and parameter estimation in case PE is large. An interesting question for future research is to extend the proposed MHE approach to achieve an estimation performance comparable to a standard MHE scheme for joint state and parameter estimation in case of large PE, while the robust stability properties irrespective of PE of the proposed approach are preserved.